OpenStudy (anonymous):

If sinΘ = 4/5 and Θ has its terminal side in Quadrant II, find the exact value of tan(2Θ).

OpenStudy (anonymous):



OpenStudy (anonymous):

tan (2theta) ?

OpenStudy (anonymous):

@surjithayer

OpenStudy (anonymous):

\[\tan ~2\theta=\frac{ 2\tan \theta }{ 1-\tan ^2 \theta },\tan \theta=-\frac{ 4 }{ 3 }\] substitute the values and simplify.

OpenStudy (anonymous):

tan 2theta = 24/7 ? @surjithayer